# Transformer Block Coupling and its Correlation with Generalization in LLMs

### v4

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we analyze the trajectories of token embeddings as they pass through transformer blocks, linearizing the system along these trajectories through their Jacobian matrices. By examining the relationships between these block Jacobians, we uncover the phenomenon of transformer block coupling in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters such as parameter count, model depth, and embedding dimension. We further investigate the emergence of these properties through training, observing the progressive development of coupling, as well as increased linearity and layer-wise exponential growth in the token trajectories. Additionally, experiments with ViTs further validate emergence of coupling and its correlation between coupling and generalization, complementing our findings in LLMs. Collectively, these insights provide a novel perspective on token interactions in transformers and open directions for studying and improving training and generalization.
@@ 1 Introduction
In our work, we investigate whether there are identifiable structural characteristics across 30+ pretrained LLMs, analyze their emergence with training, and examine their relationship with generalization performance. During inference, as token embeddings pass through the network, we analyze the their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study their relationships. To this end, we compute the Jacobians of distinct connections between layers and tokens, perform singular value decompositions (SVDs), and compare the resulting singular vectors. This approach measures the degree of coupling between singular vectors to capture the operational similarity of blocks as they act on tokens. This perspective raises several key questions:
Q1. What regularity properties do these trajectories exhibit, and what are their relations with one another? More concretely, what is the relation between the Jacobians across different tokens and transformer layers?
Q2. How do the properties of hidden representations and their relations emerge with training?
Q3. Are any of these properties related to the generalization capabilities of LLMs?

### v5

@@ Abstract
Large Language Models (LLMs) have made significant strides in natural language processing, and a precise understanding of the internal mechanisms driving their success is essential. In this work, we analyze the trajectories of token embeddings as they pass through transformer blocks, linearizing the system along these trajectories through their Jacobian matrices. By examining the relationships between these block Jacobians, we uncover the phenomenon of transformer block coupling in a multitude of LLMs, characterized by the coupling of their top singular vectors across tokens and depth. Our findings reveal that coupling positively correlates with model performance, and that this relationship is stronger than with other hyperparameters such as parameter count, model depth, and embedding dimension. We further investigate how these properties emerge during training, observing a progressive development of coupling, increased linearity, and layer-wise exponential growth in token trajectories. Additionally, experiments with Vision Transformers (ViTs) corroborate the emergence of coupling and its relationship with generalization, reinforcing our findings in LLMs. Collectively, these insights offer a novel perspective on token interactions in transformers, opening new directions for studying their mechanisms as well as improving training and generalization.
@@ 1 Introduction
In our work, we investigate whether there are identifiable structural characteristics across 30+ pretrained LLMs, analyze their emergence with training, and examine their relationship with generalization performance. During inference, as token embeddings pass through the network, we analyze their layer-wise trajectories and linearize the effect of transformer blocks on the token embeddings across model depth to study the relationships between blocks. To this end, we compute the Jacobians of distinct connections between layers and tokens, perform singular value decompositions (SVDs), and compare the resulting singular vectors. This approach measures the degree of coupling between singular vectors to capture the operational coordination of blocks as they act on tokens. This perspective raises several key questions:
Q1. What regularity properties do these trajectories individually exhibit, and how are the blocks related to each other? More concretely, what is the relationship between the block Jacobians across different tokens and transformer layers?
Q2. How do the properties of hidden representations and their interactions emerge during training?
Q3. Are any of these properties linked to the generalization capabilities of LLMs?